OpenStudy (anonymous):

Create a quadratic equation in general form and in standard form. coordinates A(0,35) which is h and k B(19,10) which is x and y y=a(x-h)2+k Solve.10=a(19-0)2+35

OpenStudy (anonymous):

10=a(19)^2+35 10-35=a(19)^2 -25=361a -25/361=a

OpenStudy (anonymous):

Thank you.